The sum of two numbers is 12. The difference between the square of the larger number and the square of the smaller number is 24. Find these two numbers.
Answer by oscargut(2103)

The sum of two numbers is 12. The difference between the square of the larger number and the square of the smaller number is 24. Find these two numbers.
Let x and y those numbers
x+y=12 (1)
x^2-y^2 = 24 (2)
From (2) (x-y)(x+y)=24
Then: (x-y)12=24
x-y = 2 (3)
Adding (1)+(3)
2x = 14
x = 7
Using (1) 7+y=12
y = 5
Answer: x = 7, y = 5
You can visit my page: http://maths2012.webs.com
Or you can send me more problems to my e-mail: mthman@gmail.com